A carnot engine operates between two reserviour at temperature T1 and T3 . The workout put of the engine is 0.6 times the heart rejected. The difference in temperature between the source and sink is 200k. calculate the thermal efficiency ,source and sink temperature


1
Expert's answer
2020-12-28T08:56:02-0500

We have two reserviour T1 and T3 .

he workout put of the engine is 0.6 times the heart rejected.

A = Q1 - Q2 = 0.6Q1

"\eta = \frac{Q_1-Q_2}{Q_1} = \frac{0.6Q_1}{Q_!} = 0.6"

T1 - T2 = 200K

"\eta = \frac{T_1 - T_2}{T_1} = \frac{200}{T_!} = 0.6" -> T1 = "\frac{2000}{6}" = 333K - source temperature

T2 = T1 - 200K = 133K - sink temperature